### v1

# Benchmarking Fake Voice Detection in the Fake Voice Generation Arms Race

Abstract
As advances in synthetic voice generation accelerate, an increasing variety of fake voice generators have emerged, producing audio that is often indistinguishable from real human speech. This evolution poses new and serious threats across sectors where audio recordings serve as critical evidence. Although fake voice detectors are also advancing, the arms race between fake voice generation and detection has become more intense and complex. In this work, we present the first large-scale, cross-domain evaluation of fake voice detectors, benchmarking 8 state-of-the-art models against datasets synthesized by 20 different fake voice generation systems. To the best of our knowledge, this is the most comprehensive cross-domain assessment conducted to date. Our study reveals substantial security vulnerabilities in current fake voice detection systems, underscoring critical gaps in their real-world robustness. To advance the field, we propose a unified and effective metric that consolidates the diverse and often inconsistent evaluation criteria previously used across different studies. This metric enables standardized, straightforward comparisons of the robustness of fake voice detectors. We conclude by offering actionable recommendations for building more resilient fake voice detection technologies, with the broader goal of reinforcing the foundations of AI security and trustworthiness.
1 Introduction
The evolution of fake voice generation has changed digital communication, allowing applications such as text-to-speech (TTS), voice conversion, and audio reconstruction to achieve unprecedented naturalness and quality [31]. Early techniques based on parametric [133, 89, 117, 118] or concatenative [77, 53, 111] methods have given way to deep neural architectures such as WaveNet [122] which introduced an autoregressive generative model for raw audio, and Tacotron [107] which employs sequence-to-sequence architectures to map text directly to spectrograms. Recently, vocoders based on generative adversarial networks (GAN) [33] and diffusion have pushed the boundaries further, achieving near-human levels of prosody and intelligibility [147, 57, 59].
This rapid evolution in synthesis quality has been applied a wide range of services, from virtual assistants and accessibility tools to customer service and content creation [109, 80]. However, as fake voice approaches human parity, it also exposes vulnerabilities in automatic speaker verification (ASV) systems and their related countermeasure (CM) systems [38, 144]. Fake voice generation poses a critical security risk in sensitive domains such as financial services, law enforcement, and biometric security [56, 50, 8]. Recent incidents have demonstrated how malicious actors can exploit high-fidelity voice conversion to impersonate executives and facilitate unauthorized transfers [1, 144].
Consequently, the arm race between advanced fake voice generation and fake voice detection has emerged [19]. Although prior studies and benchmarks have evaluated isolated components [2, 147, 116, 84, 68], few have addressed the coupled evolution of fake voice generation methods and detection systems. Prior research in fake voice detection has conventionally aggregated fake voice samples into a single homogeneous dataset for detector training and evaluation [125, 73, 141, 94, 76, 129, 47, 128, 127]. For instance, in [140], they treat detection primarily as a static evaluation task, using a fixed dataset (constructed from fake voice samples) to assess the performance of various fake voice speech detectors across standardized manipulations and scenarios. This practice ignores critical nuances: it prevents the identification of unique acoustic artifacts inherent to each generation paradigm and masks variations in detector performance across generator types. Recent studies show that acoustic models and vocoders leave unique fingerprints in fake voice waveforms [21]. These fingerprints can be used for source attribution and to improve synthesis quality.
We introduce an ecosystem-level benchmark that systematically evaluates the interplay between modern fake voice generation “engines” and detection “guardians.” Our framework challenges 20 state-of-the-art (SOTA) fake voice generation systems, including TTS, TTS with zero-shot voice conversion, and vocoder-based audio reconstructions, with 8 leading fake voice detectors using a novel one-to-one evaluation protocol. Unlike conventional assessments based on aggregated synthetic data, our approach evaluates each generator-detector pair individually to expose unique acoustic artifacts, method-specific sensitivities, and previously hidden vulnerabilities. Furthermore, by integrating explainability analysis, we clarify the underlying reasons for detector performance discrepancies, enabling the community to pinpoint the generation techniques that most threaten current defenses. Figure 1 outlines our SoK pipeline.
Figure 1: SoK pipeline for evaluating fake voice generation and fake voice detection systems which includes three parts: taxonomy of both fake voice generator and fake voice detector, cross experiments, and evaluation on both generator and detector.
Our Contributions:
  1. 1.
    Comprehensive Overview: We provide a comprehensive taxonomy of fake voice generation and fake voice detection systems, framed through a security lens. We identify critical vulnerabilities in fake voice detection systems that failing against high-fidelity fake voices. We also pinpoint the lack of explainable analysis of the reasons behind the detection behaviors, highlighting urgent security gaps.
  2. 2.
    Refine Taxonomy of Fake Voice Detection: We refined the taxonomy structure by providing generalization-optimized models in addition to the regular taxonomy of the detector pipeline, pinpointing the generalization needs, which is also demonstrating in our cross-domain evaluation.
  3. 3.
    Empirical Security Assessment and its Rationale: We conduct a large-scale, cross-domain evaluation analysis involving 20 fake voices and 8 fake voice detectors. This study not only highlights significant security vulnerabilities and quantifies performance variations across diverse generator-detector interactions, but also explains why those interaction might happen. We provide critical insights into the root causes of these vulnerabilities. Such understanding is essential for developing more robust and transparent fake voice ecosystems.
  4. 4.
    Recommended Practices for Secure Systems: We propose concrete practices for fake voice detectors and also provide guides for the development of fake voice generations that balance high fidelity with security resilience.
[…]
Paper Year Type Modeling Params Open Control Dataset Inference Zero
(M) Src Prosody Style Spk Hours Languages Speed Shot
Tacotron 2017 AM AR $\times$ $\times$ $\times$
DeepVoice 2017 AM AR $\times$ $\times$ $\times$
TransformerTTS 2018 AM Trans $\times$ $\times$ $\times$
FastSpeech 2019 AM Trans 30.1 $\times$ $\times$
Grad-TTS 2021 AM Diff 7.6 $\times$ $\times$ $\times$
VITS 2021 Fully E2E VAE $\times$
Your-TTS 2021 Fully E2E Trans+VAE $\times$
NaturalSpeech 2022 Fully E2E VAE 28.7 $\times$ $\times$ $\times$ $\times$
NaturalSpeech2 2023 NC-LM Diff 400 $\times$
VALL-E 2023 NC-LM Trans $\times$
VoiceBox 2023 AM FM 330 $\times$
NaturalSpeech3 2024 NC-LM Trans+Diff 1000 $\times$
MaskGCT-TTS 2024 NC-LM Trans+VQ-VAE 695
2 Fake Voice Generators
In this section, we provide a taxonomy of fake voice generation systems, with a specific focus on Text-to-Speech (TTS). TTS refers to the process of converting written text into natural-sounding spoken audio. TTS with zero-shot voice conversion is a technique that enables voice generation in a target speaker’s voice using only a short reference audio clip, without requiring prior training on that specific voice [15]. There are several categories of fake voices. A high-level pipeline is shown in Figure 2.
Figure 2: Fake Voice Generation Evolution. Traditional neural TTS pipelines convert text to acoustic representations, then utilize vocoders to synthesize waveforms. Fully E2E models follow a traditional approach, but optimize all components at once. Neural codec models encode and decode audio using a data-driven approach instead of static compression algorithms. Audio inputs are optional and are often used for speaker conditioning in zero-shot TTS with voice conversion.
From Figure 2, we can see two methods for arriving at the output fake voice. First, traditional approaches use a neural network-based text encoder to generate latent features based on input text. Then, a vocoder synthesizes the waveform from hidden features. Note that traditional approaches can also be end-to-end (E2E), where all components are optimized at once. Second, neural audio codecs provide a machine learning-based compression mechanism for raw audio. These approaches utilize the encoder and decoder from a neural audio codec, along with a text encoder, to synthesize raw audio. The rest of this section will expand on different methods of fake voice generation.
2.1 Traditional Generators with Vocoder
Traditional fake voice generation systems typically include two main parts: a text encoder that converts text inputs into meaningful acoustic features (the acoustic model), and a vocoder which synthesizes raw waveforms using the acoustic features.
2.1.1 Acoustic Model
Acoustic models generate acoustic representations from linguistic features or even directly from phonemes or characters [116]. Table 1 shows some of the acoustic models that we investigated, which span three main categories (more details in Table 9 of Appendix): autoregressive, flow matching & diffusion-based, and transformer-based.
Autoregressive Model: This type of acoustic model transformed fake voice generation by aggregating text-to-speech pipelines into neural networks and obtaining excellent naturalness with minimum manual feature engineering. Inspired by the Tacotron [102, 107, 134] series, these models autoregressively produce spectrograms from text using encoder-attention-decoder architectures. Later improvements included DurIAN [143] for stable alignment use duration predictors rather than attention. DeepVoice 1-3 [6, 85, 5] combined CNNs for mel spectrogram generation and linguistic feature extraction. Autoregressive systems remain fundamental despite difficulties in latency and controllability, usually coupled with non-autoregressive components such as duration predictors or equipped with advanced vocoders to balance quality and speed in contemporary systems.
Flow-Matching & Diffusion-Based Models: Flow-matching and diffusion models represent two powerful paradigms for transforming simple base distributions (e.g., Gaussians) into complex data distributions, unified under the framework of continuous-time generative modeling.
Continuous Normalizing Flows (CNFs) generalize discrete flows by formulating the transformation as the solution to an ordinary differential equation (ODE), yielding continuous-time invertible dynamics. Building on CNFs, Flow Matching techniques [71] learn velocity fields that directly transport samples from the base to the target distribution by minimizing an optimal-transport–informed loss, effectively bridging the gap between likelihood-based flows and transport-based formulations. In the realm of Text-to-Speech (TTS), models like VoiceBox [64] combine flow-matching objectives with scalable in-context learning to refine alignment trajectories. Similarly, Matcha-TTS [78] employs conditional flow matching guided by optimal transport to enhance style and prosody control.
In parallel, Diffusion Models utilize stochastic differential equations (SDEs) to gradually denoise samples, transforming them from pure noise into coherent data. A notable example in TTS is Grad-TTS [87], which leverages score-based diffusion in the spectrogram domain to achieve high-fidelity voice generation through iterative refinement.
Crucially, Flow Matching provides a principled link between deterministic Ordinary Differential Equation (ODE) flows and the stochastic SDE approaches characteristic of diffusion. By matching marginal distributions along continuous trajectories, flow matching unifies these two methodologies, combining the potential for exact density evaluation with the strong sample quality often associated with diffusion-based generative modeling.
Transformer-Based Model: The Transformer architecture [123] has revolutionized the field of text-to-speech, largely replacing earlier recurrent or convolutional approaches. Its core strength lies in the powerful self-attention mechanism within its encoder-decoder structure, enabling the effective conversion of input text sequences into compressed intermediate representations. These representations are then decoded to generate acoustic features, such as mel-spectrograms, which form the basis for fake voice generation [116]. This paradigm shift reveals in the development of numerous influential models. An early key example is Transformer-TTS [69], which successfully adapted the original Transformer by integrating it with design principles from the established Tacotron 2 architecture [102]. Seeking to overcome the latency issues inherent in autoregressive generation, the FastSpeech series emerged as a significant advancement. FastSpeech [97] pioneered non-autoregressive parallel generation for faster synthesis, relying on an external aligner to predict phoneme durations. Its successor, FastSpeech 2 [96], further refined this parallel approach by directly predicting variance information like pitch and energy within the model, simplifying the training pipeline and enhancing controllability. Subsequent research focused on improving the expressiveness and naturalness of the generated speech; DelightfulTTS [74], for instance, incorporated Conformer modules to better capture both local and global acoustic context. More recently, architectures like E2-TTS [27] explore novel directions such as flow-matching transformers operating directly on simple character sequences augmented with filler tokens, demonstrating the ongoing evolution of the Transformer acoustic model in TTS fake voice generation.
2.1.2 Vocoder
Vocoders are used for generating waveforms from acoustic features. We investigate vocoders based on two main categories: autoregressive vocoder and non-autoregressive vocoder. For latter category, we focus on three principal categories: Flow-Based, Diffusion-Based, and Generative Adversarial Network (GAN)-Based vocoders.
Autoregressive Vocoder: Unlike conventional parametric vocoders dependent on signal processing methods, autoregressive vocoders create audio samples sequentially, each new sample conditioned on prior produced samples and acoustic properties. By capturing the temporal dependencies in speech signals, Autoregressive vocoders such as WaveNet [122], WaveRNN [48], and LPCNet [121] have demonstrated remarkable capabilities in reproducing natural speech characteristics, including phase information and fine spectral details that are often lost in conventional vocoding methods. Their capacity to replicate the probabilistic character of voice waveforms has made them as foundation for contemporary neural vocoder.
Flow-Based Vocoder: Normalizing flow-based generative vocoders simplify complex probability distributions using invertible transformations, typically based on Gaussian distributions [116]. The core mechanism of these models converts data between complex and standard distributions via invertible mappings, sampling from the standard distribution and applying inverse transformations for generation. Foundation models in this approach include WaveGlow [88], which combined Glow and WaveNet for a non-autoregressive model using affine coupling layers, and WaveFlow [86], which leverages invertible transformations to enable efficient and parallel generation.
Diffusion-Based Vocoder: Diffusion vocoders leverage denoising diffusion probabilistic models [39, 116, 147] to map data and latent distributions via forward and reverse processes. The core mechanism consists of a forward process that adds noise to clean audio via a Markov chain with Gaussian transitions, and a reverse process that recovers clean audio by estimating noise with a simplified loss function. Key models in this field include foundation works such as DiffWave [58], which uses versatile Markov chain generation as the keystone diffusion-based vocoder, and WaveGrad [13], which employs gradient-based noise estimation. Noise distribution improvements have been achieved with PriorGrad [66], which introduces adaptive data-dependent priors. Several acceleration techniques have emerged, including FastDiff [41], which develops dynamic noise scheduling, and BDDM [62], which utilizes bilateral denoising diffusion modeling. Advanced optimizations continue to refine these approaches, such as FreGrad [82], which employs wavelet-based sub-band decomposition for improved efficiency. Rectified flow also has been used in diffusion models such as RFWave [72], which adapts Rectified Flow with multi-band processing to address latency issues, further improving the efficiency and quality of fake voice generation.
GAN-Based Vocoder: GAN vocoders employ adversarial training with a discriminator to assess authenticity [67, 116], enabling high-quality fake voice generation. Foundation models include MelGAN [59], which uses multi-scale discriminators and hinge loss, and ParallelWaveGAN [139], which implements multi-resolution STFT losses. Architectural improvements followed with HiFi-GAN [57] introducing multiperiod discriminators and multireceptive field fusion, UnivNet [42] featuring multiresolution discriminators for full-band signals, and BigVGAN [67] innovating with periodic activations. Artifact reduction was addressed by FA-GAN [104] with anti-aliased deconvolution and Avocado [7] mitigating spectral leakage. Digital Signal Processing integration advanced through DSP-GAN [108], SiFi-GAN [142] implementing source-filter theory, and Vocos [106] with Fourier-based time-frequency representation.
2.2 Fully End-to-End Models
Fully end-to-end text-to-speech models generate waveforms directly from text, bypassing the traditional acoustic model and vocoder pipeline with intermediate spectrograms. This approach mitigates error propagation in cascaded structures and simplifies the training process while potentially improving overall speech quality [116]. By optimizing all components in a single framework, these models can better capture the complex relationships between text and speech, resulting in higher fidelity generations. Key models in this category include the VITS series, with VITS [54] serving as a keystone model based on conditional variational autoencoder for end-to-end waveform generation, and MB-ISTFT-VITS [49] introducing multi-band generation and ISTFT decoding for enhanced speech detail. Other notable contributions include Your-TTS [10], the first work proposing a multilingual approach in the zero-shot voice conversion within multi-speaker text-to-speech scope, NaturalSpeech [115], which utilizes a differentiable durator and memory-enhanced bidirectional modeling, and Mega-TTS [43], which uses VQ-GAN with disentangled modeling for content and prosody control.
2.3 Neural Audio Codec Language Models
Neural audio codecs utilize neural networks to compress audio waveforms into compact representations via an encoder and reconstruct them using a decoder [103]. These codecs are integral components of modern text-to-speech pipelines, balancing low bit rates for compression with high audio fidelity.
Typically, these codecs generate compact representations using discrete tokens. Prominent examples like SoundStream [145], Encodec [23], and Descript Audio Codec [60] employ Vector Quantized Variational AutoEncoder (VQ-VAE) architectures often combined with residual vector quantizers for efficient encoding.
In text-to-speech, a language model predicts sequences of these discrete audio tokens based on input text, potentially conditioned on speaker embeddings or audio prompts. The codec’s decoder then generates the final audio waveform from these predicted tokens [15]. Key text-to-speech generators with zero-shot voice conversion leveraging this paradigm include VALL-E [15] which employs a two-stage approach with autoregressive and non-autoregressive transformer models as a keystone text-to-speech fake voice generator of neural codec based language model which uses Encodec [23] as tokenizer. NaturalSpeech series (2-3) [103, 44] utilizes continuous vector representations to shorten token sequences and incorporates a factorized codec design.
Recent advancements further enhance capabilities and address specific challenges and strengthen zero-shot voice conversion capabilities:
  • XTTS [9] and CosyVoice [26]: Focuses on massively multilingual support and sophisticated voice modification features.
  • MaskGCT-TTS [130]: Introduces a mask-and-predict training strategy, strengthening zero-shot voice conversion and computational efficiency.
  • CLaM-TTS [55]: Leverages probabilistic Residual Vector Quantization for superior compression in terms of token sequence length.
  • VoiceCraft [83]: Proposes a novel token rearrangement technique to improve synthesis quality.
  • FireRedTTS series [35, 36]: Integrates a semantic-aware speech tokenizer for generation from text/audio prompts and introduces a streamable codec architecture using causal convolutions for real-time applications.
The high fidelity achieved by models like NaturalSpeech3 [44] raises significant concerns about voice identity protection and potential misuse for impersonation. This necessitates research into robust detection methods and watermarking techniques for fake voice generation. Balancing rapid innovation with responsible deployment is crucial for maintaining public trust and realizing the benefits of advanced fake voice generation technologies.
3 Fake Voice Detectors
To advance the field of fake voice detection, numerous studies [68, 84] have investigated fake voice detection techniques, which can be broadly classified into two categories: pipeline and end-to-end detectors. While pipeline approaches separate feature extraction and classification, End-to-End (E2E) models operate directly on raw audio waveforms. The following section introduces both frameworks in detail.
3.1 Feature Extraction
As the front-end and a crucial component of pipeline-based detection systems, feature extraction aims to derive discriminative representations by capturing artifacts indicative of synthetic or manipulated speech. Depending on the feature generation method, existing approaches can be broadly divided into three categories: hand-crafted spectral features, deep learning (DL) features, and other analysis techniques.
3.1.1 Hand-Crafted Spectral Features
Hand-crafted features are based on human-interpretable properties of audio and leverage domain expertise to focus on characteristics known to be effective in distinguishing between real and fake voices. These features have proven to be a strong baseline in fake voice detection, offering a reliable foundation for capturing discriminative artifact patterns. Among these, spectral features describe the properties of an audio signal in the frequency domain, typically computed from the power spectrum or magnitude spectrum of the signal, which illustrate how a signal’s energy is distributed across frequencies. Notable spectral features, including the Mel Frequency Cepstral Coefficient (MFCC) [22] and the Linear Frequency Cepstral Coefficient (LFCC) [3], are particularly effective in capturing distinct patterns indicative of fake voices.
A spectrogram is a two-dimensional visual representation that illustrates how the frequency content of a signal evolves over time, typically generated by applying the Short-Time Fourier Transform (STFT) to overlapping segments of the audio signal. Among various spectrogram types, the Mel-Spectrogram and Constant-Q Transform (CQT) Spectrogram [28, 20, 92] are the most commonly used in speech processing due to their effectiveness in capturing relevant acoustic features. After extracting spectral features, supervised learning models such as Residual Networks (ResNet) and Recurrent Neural Networks (RNNs) are often employed to generate deep embeddings [51, 30]. For example, [11] utilize an U-Net structure to extract features from CQT-Spectrogram; while in [16] LFCC features and ECAPA-TDNN are combined to generate embeddings.
3.1.2 Deep Learning Features
Due to recent progress in deep learning (DL), neural network architectures are now employed to derive learnable embeddings that represent the essential traits of raw speech, complementing conventional hand-crafted features. While early approaches relied on traditional machine learning techniques and convolutional neural networks (CNNs), the field has evolved to incorporate more advanced architectures such as graph neural networks (GNNs), Transformers, and other modern deep learning models.
In recent years, self-supervised learning (SSL) frameworks have gained significant traction for feature extraction in fake voice detection, exhibiting remarkable advantages over conventional acoustic and supervised learning-based representations [75, 114]. Xie et al. [138] demonstrate the notable efficacy of Wav2Vec features when implemented within a Siamese network framework. Based on this study, Tak et al. [114] further explore the application of Wav2Vec-2.0 as a front-end feature extractor, combined with a spectro-temporal graph attention network (AASIST) as the back-end classification. Similarly, Guo et al. [37] integrate the WavLM model with a Multi-Fusion Attentive classification to enhance fake speech detection capabilities. In contrast, Wang et al. [126] employ embeddings from the pre-trained HuBERT as inputs.
3.2 Classification
Traditional back-end classifiers for fake speech detection include conventional models such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) [98, 124]. With the advancements in deep learning, convolutional neural networks (CNNs) have gained prominence for their ability to extract localized and hierarchical patterns within speech signals. Wu et al. [137] and Lavrentyeva et al. [63] implemented the Light CNN (LCNN) architecture [136] for synthetic voice identification, incorporating convolutional and max-pooling operations with Max-Feature-Map (MFM) activation to enhance discriminative feature selection.
The Residual Network (ResNet), a widely adopted CNN variant, addresses gradient vanishing issues in deep architectures through its skip-connection design. Both Alzantot et al. [4] and Tomilov et al. [119] employ ResNet as their classification component, achieving robust performance. Zhang et al. [149] introduced the Squeeze-and-Excitation Network (SENet) for fake voice detection, which enhances CNN representational capacity through channel-wise feature recalibration. ASSERT (ASqueeze-Excitation and Residual neTwork) [61], a hybrid architecture combining SENet and ResNet modules, was recognized among the top-performing systems in the ASVspoof 2019 challenge.
Graph Neural Networks (GNNs) have emerged as effective approaches in fake voice detection by modeling frequency and temporal dimensions as graph nodes [101]. Graph Convolutional Networks (GCNs) segment spectrograms into grid-like patches from which CNN-extracted embeddings are derived [12]. Tak et al. [112] proposed a Graph Attention Network (GAT) incorporating attention mechanisms to improve detection accuracy. The AASIST framework [112] enhances GAT by integrating spectral and temporal sub-graphs through a heterogeneity-aware approach.
Zhang et al. [148] introduce XLSR (Wav2Vec-2.0) with SLS Classifier, which employs a Sensitive Layer Selection module to extract discriminative features from transformer layers for fake voice detection. XLSR-Conformer [100] combines XLS-R with conformer architecture using classification tokens for variable-length utterances. Further refinement through XLSR-Conformer with TCM [120] incorporates Temporal-Channel Modeling to capture dependencies in audio signals, enhancing fake voice detection.
3.3 E2E Model
End-to-end approaches, which operate directly on raw audio waveforms without requiring explicit feature extraction, have gained significant attention in voice-related tasks, such as fake voice detection. Among these, RawNet2 [46] is a neural network originally designed for speaker verification and identification. It processes raw speech signals using a sinc convolution layer for initial feature extraction, followed by multiple residual blocks to learn deeper representations. Tak et al. [113] were the first to adapt RawNet2 for anti-spoofing tasks. Building on this, Wang et al. [131] proposed a joint optimization strategy leveraging a weighted additive angular margin loss, further improving the performance of the RawNet2-based detection system.
Another notable architecture, SincNet [91], is also designed to process raw audio directly, leveraging parameterized sinc functions to efficiently learn relevant frequency bands. Ge et al. [31] employed sinc-shaped filters and partial channel connections for efficiency. Zeinali et al. [146] introduce a fusion framework that combines SincNet and VGG architectures to enhance fake voice detection performance. Additionally, Huang et al. [40] propose DFSincNet, an end-to-end model that enhances SincNet by incorporating discriminative frequency information, with a particular focus on high-frequency components to improve spoofing detection accuracy. RawBMamba [18] utilizes sinc and convolutional layers for short-range features and bidirectional mamba for long-range features, with a fusion module integrating embeddings to enhance audio representation.
4 Cross-domain Evaluation
To address the limitations of common fake voice detector evaluations, which typically rely on static and aggregated datasets, thereby obscuring the unique artifacts introduced by different fake voice generators, we propose a cross-domain, comprehensive assessment. This evaluation includes two experiments involving 22 types of fake voices, generated using 15 distinct voice synthesis models from three major categories: traditional pipeline models, neural codec language models, and end-to-end models, along with five existing datasets for comparison. To assess the robustness and generalizability of detection systems, we evaluate these voices using 8 different fake voice detectors, analyzing their performance on both individually generated voices and aggregated datasets. Since detector performance can vary significantly depending on the type of fake voice, with some detectors excelling in specific scenarios while failing in others, our approach provides a more nuanced understanding of their vulnerabilities. Additionally, we introduce two metric systems designed to quantify the overall effectiveness of both fake voice generators and detectors from multiple perspectives, enabling fair comparison and ranking across different systems.
4.1 Experimental Design
Experiments in this study were conducted using one NVIDIA RTX 4090 GPU (24 GB VRAM), an Intel i9-10900X CPU at 3.70 GHz, 256 GB of RAM, and running Ubuntu 22.04 LTS.
Fake Voices Selection 15 state-of-the-art fake voice generators were selected covering all categories introduced in Section 2, including traditional pipeline models, codec language models, and end-to-end architectures. As traditional TTS systems involve vocoders for converting Mel-spectrograms to waveforms, we also evaluated standalone vocoder performance for audio reconstruction. To ensure comprehensive assessment, we included public datasets containing both real and fake audio: the latest ASVspoof-2024, ASVspoof-2021 LA and DF, the Fake or Real (FoR) dataset, and the CFAD dataset for bilingual analysis. A detailed overview of selected fake voices is presented in Table 2.
Table 2: Details of selected fake voices from existing studies. NC-LM=Neural Codec Language Model. *: with Zero-Shot Voice Conversion Capability
Category Type Models/Dataset
TTS Transformer-based FastSpeech2 + HiFi-GAN
YourTTS + HiFi-GAN *
Flow Matching Matcha-TTS + HiFi-GAN
Diffusion-based GRAD-TTS + HiFi-GAN
NC-LM MaskGCT-TTS *
FireRedTTS-1S *
Vocoder GAN Structure HiFi-GAN
MelGAN
ParallelWaveGAN
Arch Improved GAN Vocos
Sifi-GAN
BigVGAN
Diffusion-based Diffwave
FastDiff
RFWave
Existing
Dataset
English Dataset ASVSpoof-21LA
ASVSpoof-21DF
ASVSpoof-5
Fake or Real (FoR)
Chinese Dataset CFAD
Fake Voice Generation Process LibriSpeech is a widely used open-source dataset designed for automatic speech recognition (ASR) and text-to-speech (TTS) research. It contains English speech derived from audiobooks, paired with aligned text transcriptions. In our experiments, we used the test-clean subset, which includes 8.56 hours of audio from 39 speakers, sampled at 24 kHz in a multi-speaker English corpus. For TTS systems, we used the transcriptions as input to generate fake (synthetic) voices. The original LibriSpeech audio served as the reference for evaluating the similarity and authenticity of the generated speech. For TTS systems with zero-shot capabilities, able to mimic specific speaker characteristics, the generated fake voices closely resembled the original speaker. In contrast, TTS models without zero-shot functionality produced fake voices using a default or generic speaker profile. For vocoder evaluation (audio reconstruction), we first converted the original waveforms into Mel-spectrograms and then back into waveforms using vocoders, treating the output as synthetic speech. We compared these reconstructed outputs with the original waveforms to evaluate the vocoder’s performance. All TTS models, vocoders, and inference pipelines used in our experiments were sourced from publicly available pre-trained models on Hugging Face.
Detector Selection & Implementation Similar to the approach for selecting fake voice generators, we carefully chose 8 state-of-the-art fake voice detectors for evaluation. Our selection criteria focused on accessibility and benchmark performance on the ASVspoof dataset, which is widely used in the field for evaluating fake voice detection systems. We selected two main types of detectors: end-to-end (E2E) models and advanced models that use transformer-based front-end for feature extraction. While the latter type of detector may share a similar front-end, their back-end classifiers differ significantly, resulting in varied performance, particularly in cross-domain evaluations. The selected detectors are listed in Table 3which we divide into two categories: end-to-end detectors and detectors that used XLSR (Wav2Vec-2.0) as feature extractions. We used the official implementations of all eight detectors, each trained on the ASVspoof19-LA dataset.
Table 3: Details of selected fake voice detectors from existing studies.
Category Detection Approach
End-to-End RawNet2 [113]
RawPCDart [31]
RawBMamba [18]
AASIST [45]
Wav2Vec-2.0 Extraction XLSR-SLS [148]
XLSR-Conformer [100]
XLSR-Conformer+TCM [120]
XLSR-AASIST [114]
4.2 Evaluating Fake Voice Generators
In this section, we introduced the performance metrics used in our experiments and analyzed the correlations between them. We then proposed an overall performance score applicable to all types of generators, aimed at evaluating the effectiveness of both existing and future fake voices.
4.2.1 Performance Metric for Generator
The Mean Opinion Score (MOS) has been the standard for decades; yet, recent text-to-speech (TTS) systems often exceed the discriminative resolution of its five-point scale and suffer from listener bias, high variance, and costly test administration. [65] Consequently, we adopted a set of objective metrics for TTS and vocoder evaluation that (i) correlate with perceptual quality, (ii) cover intelligibility, naturalness, and similarity, and (iii) computational overhead. Then, we computed an overall score based on those criteria. Exact metric computation procedures are given in Appendix G.
For TTS systems (including Zero-Shot) :
PSNR [2]: Compares maximum signal power to noise power.
DNSMOS[93]: Using deep learning methods to predict subjective audio perceptual quality.
Real-time Factor (RTF): The ratio of the time taken by the model to process or generate one second of audio to the actual duration of that audio; values less than 1 indicate faster-than-real-time performance.
Parameters: The total number of trainable parameters in the model, reflecting its size, complexity, and memory requirements, measured in million.
WER & CER: Evaluate word error rate (WER) and character error rates (CER) using Whisper[90], demonstrating intelligibility.
SIM: WavLM-based [14] speaker similarity between original and generated speech pairs.
For Vocoder models:
PESQ [99]: Perceptual differences between generated and reference speech.
SSIM[132]: Frequency-domain similarity between generated and real spectrograms.
LS-MSE[2]: Mean square error between ground truth and generated log-mel spectrograms.
4.2.2 Individual Metrics Analysis and Generator Overall Performance
Each evaluation metric has its own strengths and limitations, making it suitable for specific applications. To create a more generalizable evaluation method, we introduce an overall performance score designed to assess the effectiveness of all types of voice generation systems, both current and future. Given the redundancy and complementarity among metrics used in text-to-speech (TTS) and vocoder models, we assign different weights to each metric and compute a final overall score to enable a more comprehensive performance analysis.
Generator Overall Score Formulation. We define the overall performance of a generator as a weighted aggregation of normalized metrics. For a given task, the composite score is expressed as:
$F_{\text{TTS}}$ $=\sum_{i}\alpha_{i}\cdot M_{i,\text{norm}},$ (1)
$F_{\text{Audio Reconstruction}}$ $=\sum_{j}\beta_{j}\cdot M_{j,\text{norm}},$ (2)
where $M_{i,\text{norm}}$ and $M_{j,\text{norm}}$ denote the $i$-th and $j$-th evaluation metrics, each normalized to the range $(0,1]$. The coefficients $\alpha_{i}$ and $\beta_{j}$ are non-negative weights that satisfy $\sum_{i}\alpha_{i}=1$ and $\sum_{j}\beta_{j}=1$, respectively.
The specific choice of metrics (DNSMOS, SIM, PESQ, PSNR, SSIM, WER, RTF, and model size) and their associated weights are detailed in Appendix G.
Table 4: Performance evaluation of various TTS generators. Column best values highlighted blue, worst red. WER and CER are shown in %. Params = Parameters in Million, Overall = Overall score.
Method DNSMOS $\uparrow$ WER $\downarrow$ CER $\downarrow$ PSNR $\uparrow$ SIM $\uparrow$ RTF $\downarrow$ Params (M) Overall $\uparrow$
FastSpeech2 + HiFi-GAN 3.1140 18.41 8.52 20.4370 0.6650 0.2403 71 0.291
GRADTTS + HiFi-GAN 2.9018 21.41 9.05 22.5353 0.7147 0.0658 29 0.307
MatchaTTS + HiFi-GAN 3.3067 14.94 4.85 20.4433 0.7099 0.0871 21 0.586
YourTTS (zero-shot) 3.2939 15.93 7.97 22.8675 0.9189 0.2368 87 0.829
MaskGCT-TTS (zero-shot) 3.2675 14.53 5.79 21.8595 0.9567 0.3850 2226 0.734
FireRedTTS-1S (zero-shot) 3.2727 16.10 7.53 21.6621 0.9533 0.1768 1294 0.735
Ground Truth 3.3414 N/A N/A N/A N/A N/A N/A N/A
Table 5: Performance evaluation of various Vocoders (audio reconstruction). Column best values highlighted blue, worst red. Params = Parameters in Million, Overall = Overall score.
Method DNSMOS $\uparrow$ PESQ $\uparrow$ LS-MSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RTF $\downarrow$ Params (M) Overall $\uparrow$
ParallelWaveGAN 3.0895 2.8475 0.0012 25.2606 0.8478 0.0094 17 0.4541
MelGAN 3.0259 2.9971 0.0007 26.7588 0.8831 0.0013 21 0.4716
HiFi-GAN 3.0372 3.1048 0.0006 26.7875 0.9418 0.0067 14 0.5607
DiffWave 3.1415 2.3865 0.0068 21.6858 0.7609 0.1152 3 0.2473
FastDiff 3.1211 3.2667 0.0093 19.4235 0.7550 0.0439 15 0.3485
SiFi-GAN 3.0416 3.1653 0.0009 25.8332 0.9551 0.0081 11 0.5747
Vocos 3.2325 3.5996 0.0008 27.1873 0.9751 0.0705 14 0.7890
BigVGAN 3.2797 4.0125 0.0007 28.3506 0.9812 0.0269 112 0.9081
RFWave 3.0553 3.8183 0.0089 19.7259 0.7598 0.0522 18 0.3495
Ground Truth 3.3414 N/A N/A N/A N/A N/A N/A N/A
Individual Metrics. Table 4 shows the performance of TTS and TTS with zero-shot voice conversion methods, while Table 5 shows the detailed performance evaluation of audio reconstruction methods. For audio reconstruction, BigVGAN’s strong performance across multiple perceptual (DNSMOS, PESQ) and signal-level metrics (PSNR) suggests a robust architecture capable of generating both realistic and accurate waveforms, potentially making its output more challenging for fake voice detection systems. Conversely, FastDiff’s low PSNR and MelGAN’s lower DNSMOS might reflect trade-offs that create detectable signatures for detection systems.
For TTS generators, Matcha-TTS achieving top DNSMOS suggests it excels in overall naturalness, raising the difficulty for detectors. For TTS with zero-shot voice conversion, MaskGCT-TTS and FireRedTTS-1S’s top-level scoring in multiple metrics (e.g., SIM and PSNR) reflects their strong voice conversion ability and the perceptual fidelity that neural codec provides. YourTTS, as representative of end-to-end TTS with zero-shot conversion, achieves voice conversion with fewer parameters than neural codec language model TTS, supporting lightweight computational efficiency.
Overall Performance Results. Flow Matching approaches (MatchaTTS + HiFi-GAN) perform commendably with good DNSMOS scores. For ZSVC systems, YourTTS stands out with high overall score and competitive PSNR/SIM performance while maintaining moderate parameter count (87M). Neural Codec models like FireRedTTS-1S and MaskGCT-TTS deliver strong results but with substantially larger parameter counts (1294M and 2226M respectively). For vocoders, architecture-improved GANs such as BigVGAN and Vocos demonstrate leading performance, while diffusion-based models generally exhibit lower overall scores despite having fewer parameters.
4.3 Cross-Domain Evaluation Performance Analysis
We perform a comprehensive analysis based on Table Table 6 and 7 with three focus. In Section 4.3.2, we discuss one generator performance across detector. In Section 4.3.3, we discuss one detector performance across different generator. Then finally, we provide our overall score analysis in Section 4.3.4 from the both ways to evaluate the overall detector performance.
Table 6: Performance Individual metrics (EER(%), $C_{llr}$, minDCF) - Part 1. Column best (lowest) values are blue, worst (highest) red. All metrics are the lower the better.
Detector Matcha‑TTS FastSpeech2 GRAD‑TTS YourTTS MaskGCT‑TTS
EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF
RawNet2 47.88 4.967 0.997 23.10 2.395 0.873 38.93 4.970 0.922 30.31 4.955 0.867 62.44 5.198 0.999
RawPCDart 48.36 0.998 0.998 33.08 0.991 0.915 49.97 1.013 0.999 49.46 0.998 0.988 54.81 1.025 1.000
AASIST 41.70 1.639 0.998 14.24 2.393 0.843 37.69 1.614 0.939 42.67 1.673 0.982 56.90 5.134 0.998
RawBMamba 24.29 2.330 0.688 49.16 4.383 0.998 23.07 2.212 0.604 34.22 3.166 0.891 65.18 6.796 0.999
XLSR‑AASIST 35.85 2.259 0.868 41.22 2.242 0.922 3.42 2.397 0.073 28.57 1.662 0.587 45.19 2.573 0.977
XLSR‑Conformer 49.85 1.729 0.995 26.72 1.695 0.764 2.31 1.672 0.066 15.49 1.682 0.419 41.85 1.732 0.906
XLSR‑SLS 48.36 3.095 0.998 2.58 2.910 0.070 44.34 3.065 0.996 5.46 2.910 0.149 24.94 2.951 0.626
XLSR‑Conf.+TCM 43.89 1.576 0.911 30.34 1.298 0.852 3.49 1.232 0.100 9.59 1.244 0.268 27.00 1.333 0.677
Detector FireRedTTS‑1S RFWave SiFiGAN BigVGAN HiFi‑GAN
EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF
RawNet2 60.25 5.095 1.000 32.57 2.110 0.797 25.22 2.057 0.613 43.69 2.218 0.997 46.07 2.239 0.998
RawPCDart 53.94 1.327 1.000 44.69 0.991 1.000 44.40 0.964 0.998 44.44 0.974 0.999 44.31 0.984 0.998
AASIST 57.88 5.573 1.000 43.66 1.484 0.946 47.54 1.709 0.989 45.13 1.533 0.999 40.59 1.333 0.947
RawBMamba 71.25 6.805 1.000 49.33 2.034 0.952 44.82 1.626 0.925 48.60 2.034 0.977 41.15 1.492 0.923
XLSR‑AASIST 49.15 3.651 0.990 39.42 1.605 0.912 35.88 1.362 0.845 42.59 1.809 0.997 27.12 0.977 0.573
XLSR‑Conformer 46.57 2.036 0.964 37.65 1.012 0.892 30.42 0.951 0.769 32.24 0.967 0.802 11.41 0.809 0.269
XLSR‑SLS 36.59 3.362 0.737 30.94 1.153 0.729 14.37 0.991 0.376 25.92 1.111 0.648 5.16 0.925 0.136
XLSR‑Conf.+TCM 30.86 1.879 0.712 23.58 0.764 0.529 20.47 0.695 0.529 30.36 0.974 0.719 6.44 0.400 0.161
4.3.1 Performance Metric for Fake Voice Detectors
Following the ASVSpoof Challenge methodology (e.g., [125]), we adopted the minimum normalized Detection Cost Function (minDCF) and Equal Error Rate (EER) as primary metrics for evaluating countermeasure systems. Additionally, the Log Likelihood Ratio Cost ($C_{llr}$) is often used to evaluate not only discrimination but also the calibration of detector scores. To emphasize the overall fake voice detector performance across various generators while also considering model efficiency, we propose a novel Detector Overall Score ($S_{ovrl}$).
We computed a single unified overall score ($S_{i}$) for each fake voice detector $i$. This score combines the detector’s empirical performance against multiple fake voice generators with a penalty for its model size. The construction is as follows:
First, raw performance metrics for detector $i$ against generator jj—EER ($m_{ij}^{(\text{EER})}$), $C_{llr}$ ($m_{ij}^{(C_{\text{llr}})}$), and minDCF ($m_{ij}^{(\text{minDCF})}$), generically denoted $m_{ij}$—were linearly rescaled to $[0,1]$. This normalization uses the global minimum ($m_{\text{min}}$) and maximum ($m_{\text{max}}$) values for that specific metric type:
$n_{ij}=\begin{cases}\frac{m_{ij}-m_{\text{min}}}{m_{\text{max}}-m_{\text{min}}},&\text{if }m_{\text{max}}\neq m_{\text{min}},\\ 0,&\text{otherwise.}\end{cases}$
This yields normalized metrics $n_{ij}^{(\text{EER})}$, $n_{ij}^{(C_{\text{llr}})}$, and $n_{ij}^{(\text{minDCF})}$, where lower values indicate better performance.
Next, to emphasize challenging fake voice generation methods (e.g., Text-to-Speech (TTS), TTS with Voice Conversion, Audio Reconstruction), their contributions were weighted. For each generator $j$, a challenge score $R_{j}$ was assigned, with higher $R_{j}$ for more challenging generators. These $R_{j}$ scores were normalized to produce weights $w_{j}$:
$w_{j}=\frac{R_{j}}{\sum_{k}R_{k}}$
where the sum is over all generator challenge scores. For example, if for three generators $G_{1},G_{2},G_{3}$, their assigned scores are $R_{1}=2,R_{2}=1,R_{3}=1$ respectively (total $\sum_{k}R_{k}=4$), the weights become $w_{1}=0.5,w_{2}=0.25,w_{3}=0.25$. This ensures more difficult generators receive greater impact.
The average score $A_{ij}$, detector $i$’s combined performance against generator $j$, is the mean of its three normalized metrics:
$A_{ij}=\frac{1}{3}\left(n_{ij}^{(\text{EER})}+n_{ij}^{(C_{\text{llr}})}+n_{ij}^{(\text{minDCF})}\right)$
Overall empirical performance $P_{i}$ for detector $i$ is the weighted sum of $A_{ij}$ scores over all generators $j$:
$P_{i}=\sum_{j}w_{j}A_{ij}$
For model complexity, parameter counts $C_{i}$ were log-transformed ($\log C_{i}$) and normalized. Let $L_{\text{min}}=\min_{k}\log C_{k}$ and $L_{\text{max}}=\max_{k}\log C_{k}$. The normalized log-parameter count $\hat{C}_{i}$ is:
$\hat{C}_{i}=\begin{cases}\frac{\log C_{i}-L_{\text{min}}}{L_{\text{max}}-L_{\text{min}}},&\text{if }L_{\text{max}}\neq L_{\text{min}},\\ 0,&\text{otherwise.}\end{cases}$
This maps $\log C_{i}$ to $\hat{C}_{i}\in[0,1]$, where higher values mean larger, less efficient models.
Finally, the overall score $S_{i}$ blends empirical performance $P_{i}$ and model complexity penalty $\hat{C}_{i}$ using a trade-off factor $\alpha\in[0,1]$:
$S_{i}=\alpha P_{i}+(1-\alpha)\hat{C}_{i}$
The resulting $S_{i}$ is a consolidated, lower-is-better score, reflecting both detector accuracy and efficiency.
4.3.2 Generalization Analysis of Fake Voice Generators
This section analyzes how state-of-the-art detectors generalize across different families of fake voice generators. We discuss performance trends using the Equal Error Rate (EER, %) metrics reported in Table 6 and 7.
Aggregated Datasets. While detectors perform well on standard benchmarks like ASVSpoof-21LA, with top systems achieving EERs below $2\%$ (e.g., $0.82\%$ for XLSR-AASIST), their performance degrades significantly under distribution shifts. EERs increase on compressed audio like ASVSpoof-21DF (e.g., XLSR-AASIST to $2.85\%$) and data with waveform degradations in FoR (XLSR-SLS at $4.07\%$). The challenge is most severe in cross-lingual settings, such as the Chinese CFAD dataset, where even robust models like XLSR-SLS see EERs rise to $13.06\%$, and weaker ones like AASIST approach chance-level performance ($49.11\%$). This highlights a critical weakness in out-of-distribution generalization.
Audio Reconstruction: GANs and Diffusion Models. Detectability of audio reconstruction models correlates strongly with architectural maturity.
  • GAN-based Models: Early GANs like MelGAN, HiFi-GAN, and ParallelWaveGAN are easily identified by exploiting artifacts like phase misalignment, with a top EER of $0.57\%$ (XLSR-SLS). However, advanced architectures that mitigate these flaws are far more challenging. Best-case EERs rise to $14.37\%$ for SiFiGAN, $17.49\%$ for Vocos, and $25.92\%$ for BigVGAN.
  • Diffusion-based Models: A similar trend holds for diffusion vocoders. Early models like DiffWave and FastDiff leave residual high-frequency noise, making them detectable with low EERs (e.g., $1.00\%$ for XLSR-SLS against DiffWave). In contrast, newer variants like RFWave, which better suppress such artifacts, are significantly harder to detect, pushing the best EER to $23.58\%$.
Table 7: Performance Individual metrics (EER(%), $C_{llr}$, minDCF) - Part 2. Column best (lowest) values are blue, worst (highest) red. All metrics are the lower the better.
Detector MelGAN DiffWave FastDiff Parallel‑WaveGAN Vocos
EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF
RawNet2 39.21 2.174 0.933 41.92 2.206 0.985 26.96 2.061 0.695 40.52 2.191 0.936 43.71 2.219 0.951
RawPCDart 43.45 0.976 0.996 44.79 1.000 1.000 42.30 0.950 0.924 43.69 0.979 0.996 51.90 1.029 1.000
AASIST 22.10 0.682 0.618 34.90 1.133 0.917 36.85 1.215 0.821 29.86 0.939 0.806 44.77 1.522 0.953
RawBMamba 35.17 1.203 0.856 47.23 1.488 0.999 34.88 1.528 0.886 29.98 1.817 0.746 48.13 2.903 0.997
XLSR‑AASIST 7.64 0.425 0.144 17.54 0.572 0.372 24.72 0.830 0.555 13.25 0.484 0.257 40.64 1.698 0.915
XLSR‑Conformer 0.77 0.764 0.019 3.32 0.773 0.080 15.20 0.832 0.370 1.47 0.767 0.038 25.88 0.914 0.649
XLSR‑SLS 0.57 0.914 0.016 1.00 0.914 0.031 4.51 0.924 0.118 0.90 0.914 0.026 17.49 1.018 0.429
XLSR‑Conf.+TCM 0.74 0.352 0.021 1.49 0.353 0.040 8.06 0.422 0.198 10.13 1.015 0.286 20.87 0.705 0.499
Detector FoR CFAD ASVSpoof‑5 ASVSpoof‑21LA ASVSpoof‑21DF
EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF EER $C_{llr}$ minDCF
RawNet2 17.77 4.309 0.507 43.20 1.699 0.934 36.04 4.094 0.827 9.50 1.274 0.258 22.38 3.154 0.625
RawPCDart 43.26 0.959 0.935 46.76 1.035 0.990 46.22 1.008 1.000 13.72 0.798 0.375 24.84 0.907 0.908
AASIST 12.32 1.084 0.312 49.11 1.356 0.997 29.12 4.001 0.711 11.46 0.851 0.329 26.75 4.440 0.616
RawBMamba 24.49 1.584 0.528 34.78 2.324 0.850 38.68 2.612 0.884 3.19 0.435 0.125 15.85 2.951 0.426
XLSR‑AASIST 6.83 1.313 0.160 16.03 0.670 0.452 6.77 1.231 0.136 0.82 0.054 0.023 2.85 0.081 0.598
XLSR‑Conformer 4.40 1.199 0.120 15.69 1.158 0.427 1.98 1.039 0.049 1.88 0.458 0.054 3.09 0.543 0.082
XLSR‑SLS 4.07 1.349 0.116 13.06 2.889 0.337 1.73 1.247 0.048 2.87 1.199 0.078 1.92 1.485 0.050
XLSR‑Conf.+TCM 5.01 0.603 0.141 14.69 0.953 0.388 1.64 0.473 0.047 1.03 0.202 0.056 2.06 0.504 0.089
Neural Codec Language Model TTS This family, including MaskGCT-TTS and FireRedTTS-1S, represents the most difficult detection scenario. These models leverage a neural audio codec as a final polishing stage, explicitly optimizing for perceptual realism. Their advanced techniques—such as mask-predict paradigms, two-step decoding, and zero-shot voice conversion—produce speech with fewer artifacts and more natural prosody. Consequently, detector performance is severely compromised. The best EER against MaskGCT-TTS is $24.94\%$ (XLSR-SLS), and it rises to $30.86\%$ for the even more advanced FireRedTTS-1S (XLSR-Conf.+TCM). For these models, many detectors fail entirely, with EERs exceeding $60\%$.
4.3.3 Generalization Analysis of Fake Voice Detectors
We evaluate detector performance across two main categories, analyzing their generalization trends, underlying causes, and inherent trade-offs.
End-to-End Models End-to-end models generally exhibit limited generalization, characterized by high error rates and poor calibration when faced with diverse and unseen speech generators. Performance within this category is inconsistent; for instance, RawNet2 often struggles across most generators, while RawPCDart can excel against specific types. A common weakness is unstable calibration, with models like RawNet2 and RawBMamba exhibiting high $C_{llr}$ scores.
This brittleness arises because standard end-to-end training objectives (e.g., cross-entropy) do not explicitly enforce robustness to domain shifts. The convolutional front-ends used in models like RawNet2 [113] and RawBMamba [18] learn features that are highly discriminative for in-domain data but fail to generalize to novel synthesis artifacts. Even AASIST [45], despite its improved graph attention classifier, is ultimately constrained by its RawNet2-based encoder and shows more generalization failures than detectors with more specialized architectures.
Wav2Vec-2.0 Feature Extratcion Models This category of models demonstrates highly context-dependent performance, achieving state-of-the-art results against certain generator families but failing significantly on others. For example, XLSR-AASIST excels on in-domain data like ASVSpoof-21LA ($0.82\%$ EER) but is less effective against modern TTS systems. This sensitivity to generator type is a defining characteristic: performance often collapses against advanced unseen systems (e.g., Matcha-TTS, FireRedTTS-1S) while remaining strong against older GAN or diffusion models. Among these models, XLSR-Conformer+TCM often demonstrates superior score calibration.
This behavior is rooted in strong architectural inductive biases that lead to specialization. The learned layer weighting in XLSR-SLS [148] or the specific interplay of local and global features in XLSR-Conformer [100] become highly optimized for the artifact types seen during training. While this specialization is effective for known spectral or temporal discontinuities, it can render the models blind to the subtle, high-fidelity artifacts or unnatural prosody in advanced TTS outputs.
Takeaways. No single detector achieves universal robustness. Feature-engineered models excel against certain generators but fail on novel ones, while end-to-end models show broader but weaker generalization.
4.3.4 Detector Overall Performance
The discrepancy between EER, $C_{llr}$, and minDCF highlights the importance of considering multiple metrics for a complete picture. Cases like XLSR-SLS (low EER, high $C_{llr}$ in CFAD evaluation) show good separation but unreliable scores, while RawPCDart (sometimes low $C_{llr}$, high EER/minDCF) shows potentially calibrated but highly indiscriminative separation. We also find generally audio reconstruction has better calibration performance than those of TTS synthesis possibly because of the mechanisms are directly reconstructing audio features which leads to better‐calibrated confident estimates. Shih et al. [105] suggest detectors are securing onto specific, potentially unreliable cues. When these cues are altered or masked [105], the detector’s confidence/scoring mechanism likely breaks down, which would manifest as poor calibration. Those scenarios highlight limitations not captured by single metric alone, demonstrating the necessity of multiple metrics evaluations.
Table 8: Parameter count and Overall scores ($S_{i}$) for various fake voice detectors. Lower scores indicate better overall performance. Bold indicates the best value in each column.
Detector Params (M) Overall Score ($S_{i}$) $\downarrow$
RawPCDart 24.50 0.7739
RawNet2 25.43 0.7030
RawBMamba 0.72 0.5681
XLSR-AASIST 317.84 0.5073
AASIST 0.30 0.5053
XLSR-Conformer 319.74 0.4239
XLSR-SLS 340.79 0.3870
XLSR-Conf.+TCM 319.77 0.3855
The overall detector scores ($S_{i}$) presented in Table 8 synthesize empirical accuracy ($P_{i}$) and model size penalty ($\hat{C}_{i}$), building upon the detailed generator challenges discussed in Section 4.3.2 and detector-specific traits from Section 4.3.3. With empirical performance heavily weighted ($\alpha=0.8$) and EER/minDCF prioritized within $P_{i}$, lower $S_{i}$ scores indicate superior overall standing. The XLSR-based models, exemplified by XLSR-Conformer+TCM ($S_{i}=0.3855$) and XLSR-SLS ($S_{i}=0.3870$), achieved the top ranks. Their excellent average $P_{i}$ values—reflecting strengths such as good calibration and effectiveness against many generator types—successfully offset their large model sizes (high $\hat{C}_{i}$), despite known challenges with the most advanced TTS systems (e.g., MaskGCT-TTS, FireRedTTS-1S) highlighted in earlier sections.
Conversely, very lightweight models like AASIST ($S_{i}=0.5053$) and RawBMamba ($S_{i}=0.5681$) attained competitive mid-tier scores. Their minimal $\hat{C}_{i}$ penalties compensated for what Section 4.3.3 described as more modest or variable empirical performance (e.g., "limited generalization" or higher $C_{llr}$ for some end-to-end models). Detectors such as RawPCDart ($S_{i}=0.7739$) and RawNet2 ($S_{i}=0.7030$) ranked lower, as their less competitive $P_{i}$ scores were not sufficiently counterbalanced by their model complexities. These $S_{i}$ rankings thus provide a holistic but weight-dependent view, confirming that while the newest generator architectures pose significant detection hurdles (in Section 4.3.2), detectors excelling in broad empirical accuracy can lead when model efficiency is a secondary but still influential consideration.
5 Future Directions
In this section, we outline concrete and forward-looking research avenues to navigate the escalating arms race between synthetic voice generation and detection, highlighting needs of both technical innovations and systemic strategies.
  1. 1.
    Generator Design with Holistic Risk Assessment. Future research in voice generation will likely be dominated by powerful architectures like Neural Codec Language Models and end-to-end TTS. We advocate for a holistic evaluation framework that assesses not only model performance but also inherent security risks. Progress should be measured along four critical axes: (i) perceptual quality, (ii) signal and spectrum fidelity, (iii) computational efficiency, and (iv) biometric security risk (i.e., spoofing capability). Furthermore, responsible development must include a forensic analysis of potential artifacts to proactively inform the detection community.
  2. 2.
    Advancing Detector Robustness. To counter sophisticated generators, detector research must prioritize generalization and reliability. We recommend focusing on three key areas:
    • Hybrid and Ensemble Designs. Fuse powerful Self-Supervised Learning (SSL) front-ends (e.g., XLSR) with diverse, specialized back-ends. Such ensembles can capture complementary acoustic and spectral cues, leading to more robust decisions than single architecture.
    • Principled Generalization Techniques. Incorporate methods to improve out-of-distribution performance. Techniques such as domain generalization, adversarial training, few-shot learning, and disentanglement losses [17, 95] are critical for mitigating the sharp performance drops observed on unseen generators.
    • Reliable Score Calibration. Develop detectors whose likelihood ratio scores remain well-calibrated under distribution shifts. To be practically useful in real-world applications, detectors must close the gap between achieving a low EER and a high decision cost ($C_{llr}$).
6 Conclusion
Our systematic evaluation shows that state-of-the-art fake voice generators consistently bypass today’s leading fake voice detectors, highlighting a fundamental trade-off between fidelity and security. At the same time, no single detection method proved robust against every fake voice generation techniques we tested, underlining serious gaps in current fake voice detection practices. To secure the future voice ecosystem, fake voice generator designs could embed forensic-aware constraints to be more secured while detector research must prioritize adaptability and cross-method generalization to stay ahead of a rapidly diversifying threat landscape.
Appendix A Ethical Considerations
The development and evaluation of fake voice generation and detection technologies raise critical ethical concerns. While progress in synthetic voice technologies supports beneficial applications such as accessibility for speech-impaired individuals, educational tools, and creative content production, it simultaneously creates risks for fraud, impersonation, misinformation, and erosion of social trust. Our research acknowledges this dual-use nature and adopts a framework grounded in The Menlo Report principles—Beneficence, Respect for Persons, Justice, and Respect for Law and Public Interest.
Stakeholders.
Key stakeholders include (i) the general public, who may benefit from stronger detection systems but could be harmed by misuse of synthetic voices; (ii) companies and organizations that rely on secure authentication and communication systems; (iii) the research community, which benefits from reproducibility and benchmarking; and (iv) the research team itself, which must ensure its work does not contribute to misuse.
Harms and Risks.
Two broad categories of potential harms are considered: tangible harms and violations of rights. Tangible harms include financial loss, reputational damage, or security breaches stemming from voice spoofing. Rights-based harms include the violation of privacy expectations, unauthorized use of personal voice likeness, and erosion of informed consent in data collection. Additionally, synthetic voices could exacerbate social harms such as misinformation campaigns or targeted harassment.
Mitigations.
We proactively mitigate these risks by: (1) restricting experimentation to publicly available datasets and pre-trained models; (2) avoiding the release of maliciously usable or identity-targeted audio; (3) limiting all generated content strictly to controlled benchmarking and analysis; (4) not deploying generated audio in real-world systems; and (5) encouraging safeguards such as controlled access, watermarking, and forensic traceability in downstream use. Risks that cannot be fully eliminated, such as the potential misuse of broader methodological insights, are acknowledged and documented transparently.
Decision to Proceed.
We judged that the ethical benefits—improving the robustness of fake voice detection and highlighting system vulnerabilities—outweigh the risks. This aligns with the Beneficence principle by maximizing potential societal benefit while minimizing foreseeable harm. Our analysis also respects the rights of individuals by not using private or sensitive data and by ensuring research outcomes are oriented toward defensive, rather than offensive, applications.
Responsible Disclosure and Community Norms.
Where vulnerabilities in detection systems are identified, we follow responsible disclosure practices, providing advance notice to relevant stakeholders (e.g., detection model developers) before public release of findings. This approach reduces potential negative outcomes while ensuring that security improvements are actionable.
Researcher Wellbeing.
We also acknowledge that exposure to adversarial audio datasets can impose cognitive or psychological strain on research team members. Appropriate measures, such as workload distribution and technical filters, were implemented to mitigate these risks.
Conclusion.
Our work is conducted with a commitment to both ethical integrity and legal compliance. While no research of this nature is without risk, our structured approach to stakeholder analysis, harm mitigation, and responsible dissemination provides a defensible ethical justification for both the conduct of this research and its publication. We encourage future researchers to build upon our work with equal or greater attention to ethical responsibility.
Appendix B Open Science
In order to advance transparency, reproducibility, and community-driven research, we are committed to open science practices. Due to the large size of the benchmark dataset used in our cross-domain evaluation, we provide a partial release at the following link: https://1drv.ms/u/c/e21f0906cf1459ef/Ef8uGy3nQdNFhYBBGKBQlFQBiOCOfoYFrN3JK6-z0iGy6A?e=6dccQU. We also have evaluation code in https://anonymous.4open.science/r/benchmark_voice-256F/ for fake voice detection which forked from ASVSpoof-5 we used in cross evaluation (including our composite generation and detection metric evaluation). Upon publication, we will release the complete and fully curated dataset and code to the community to ensure unrestricted access and facilitate future research.
Appendix C Related Work
Surveys and Benchmarks of Fake Voice Generation & Detection. Tan et al. [116] and Zhang et al. [147] discuss the rapid progress in neural TTS and diffusion models, respectively. For fake voice detection, Khan et al. [52] and Li et al. [68] offer taxonomies and pipeline analyses up to early 2024. Literature such as VocBench [2] and VoiceWukong [140] provide standardized benchmarking, while CodecFake+ broadens the corpus to unseen neural codec attacks and shows that diverse codec training improves robustness [17].
Generalization and Cross Evaluation. Performance often collapses when detectors face unseen generators or corpora [81, 105, 34, 32]. Doan et al. [24] identify highly transferable artifacts from large E2E-GAN generators, and Mishra et al. [79] use Shapley-based explanations to pinpoint shortcut features. Wu et al.’s CodecFake [135] confirms that retraining on neural codec data enhances cross-dataset resilience.
Appendix D Fake Voice Generator
Table 9: An overview of additional TTS models, supplementing Table 1. Modeling: Trans=Transformer, Diff=Diffusion, AR=Autoregressive, VAE=Variational Autoencoder, VQ=Vector Quantization, GAN=Generative Adversarial Network, FM=Flow Matching. Type: AM=Acoustic Model, Fully E2E=Fully End-to-End, NC-LM=Neural Codec Language Model. Params (M): Model parameters in millions. Open Src: ✓=Open Source, $\times$=Not Open Source. Control (Prosody/Style): ✓=Supported, $\times$=Not Supported. Dataset Spk: =Single Speaker, =Multi-Speaker (2-1000), =Large Multi-Speaker (1000+). Dataset Hours: =<50 Hours, =50-100 Hours, =>100 Hours. Dataset Languages: =Single Language, =Few Languages (2), =Multilingual (3+). Inference Speed: =Very fast (10x+ real-time), =Fast (2-10x), =Real-time or slower. Zero/One Shot: ✓=Supports zero-shot, $\times$=No voice conversion. Values are based on original papers. – indicates value not stated.
Paper Year Type Modeling Params Open Control Dataset Inference Zero
(M) Src Prosody Style Spk Hours Languages Speed Shot
Delightful-TTS 2021 AM Trans $\times$ $\times$ $\times$
HierSpeech++ 2022 NC-LM VAE 97
Matcha-TTS 2023 AM FM 18.2 $\times$ $\times$ $\times$
Mega-TTS 2023 Fully E2E VQ-GAN 222.5 $\times$
UniCATS 2023 NC-LM VQ-Diff+Trans $\times$
VoiceCraft 2024 NC-LM Trans 830
E2-TTS 2024 AM Trans + FM 335
CLaM-TTS 2024 NC-LM VAE $\times$
FireRedTTS 2024 NC-LM Trans 400
X-TTS 2024 NC-LM VQ-VAE+Trans 482
CosyVoice 2024 NC-LM FM+Trans
Table 9 provides an extended overview of recent TTS models beyond those in the main table, highlighting their modeling approaches (e.g., Transformer, Diffusion, VAE, GAN, Flow Matching), type (acoustic, end-to-end, or neural codec LM), parameter sizes, availability of open-source implementations, prosody/style control, dataset scale in terms of speakers, hours, and languages, as well as inference speed and zero/one-shot capability. This supplementary comparison illustrates the diversity of design choices in fake voice generation systems.
Table 10: Metrics, normalization range, and aggregation weights.
Task Metric Range Weight
TTS DNSMOS $(0,1]$ $\alpha_{1}=0.25$
SIM $(0,1]$ $\alpha_{2}=0.25$
PSNR $(0,1]$ $\alpha_{3}=0.15$
WER $(0,1]$ $\alpha_{4}=0.20$
RTF $(0,1]$ $\alpha_{5}=0.075$
LogParams $(0,1]$ $\alpha_{6}=0.075$
Audio Recon. DNSMOS $(0,1]$ $\beta_{1}=0.25$
PESQ $(0,1]$ $\beta_{2}=0.225$
PSNR $(0,1]$ $\beta_{3}=0.15$
SSIM $(0,1]$ $\beta_{4}=0.225$
RTF $(0,1]$ $\beta_{5}=0.075$
LogParams $(0,1]$ $\beta_{6}=0.075$
Appendix E Fake Voice Datasets
Fake voice datasets are the cornerstone of developing and evaluating fake voice detection systems, and the chosen algorithms used to generate the fake voices in the dataset define its quality. The composition and characteristics of the real and generated data profoundly influence detector design, performance, and ultimately, their generalization capabilities. Table 11 presents an overview of commonly used and recent novel datasets for training, development and evaluation. We also note the generator types associated with each fake voice dataset.
Table 11: Overview of fake voice detection datasets. Audio condition: Type of audio degradation present (clean, noisy or codec). #spoof:#bona: Approximate number of spoofed vs. bonafide utterances (k=thousands, M=millions). Unseen attack: ✓=Contains attacks/generators potentially unseen during system training phase, $\times$=All attacks/generators are known during training evaluation setup. Language: EN=English, ZH=Chinese, JP=Japanese, Multiple=more than one language. Generator type: Synthesis or transduction methods used for creating spoofing attacks (✓=Present). Sub-columns: DSP=Digital Signal Processing, AR=Autoregressive models, GAN=Generative Adversarial Network-based, Diff=Diffusion-based, Flow=Flow-based, NCodec=Neural Codec based, F-E2E=Fully End-to-End models, Trans=Transformer-based. Empty entry means not specified or not mentioned in the dataset paper.
Dataset Year Condition #spoof:#bona Unseen Language Generator type
attack DSP AR GAN Diff Flow NCodec F-E2E Trans
ASVspoof-19LA[129] 2019 Clean 90k:10k EN
FoR [94] 2019 Clean 87k:108k EN
ASVspoof-21LA[73] 2021 Codec 133k:15k EN
ASVspoof-21DF[73] 2021 Codec 520k:15k EN
Wavefake [29] 2021 Clean 118k:0 $\times$ EN, JP
Into The Wild [81] 2022 Noisy 12k:20k EN
CFAD [76] 2023 Noisy, Codec 77k:39k ZH
LibriSeVoc [110] 2023 Clean 79k:13k $\times$ EN
DFADD [25] 2024 Clean 164k:44k $\times$ EN
CVoiceFake [70] 2024 Codec 1.25M:0 $\times$ Multiple
ASVspoof-5 [125] 2024 Noisy, Codec 288k:1.21M EN
CodecFake [135] 2024 Codec 44k:44k $\times$ EN
CodecFake+ [17] 2025 Codec 90k:1.42M EN
Pivotal datasets in the field, particularly those from the ASVspoof challenge series (ASVspoof19, ASVSpoof-21, ASVspoof-5) [129, 73, 125], have been instrumental in standardizing evaluation protocols and driving progress. These challenges typically define specific tasks (e.g., Logical Access focusing on TTS and voice conversion and Physical Access on replay) and provide curated training, development, and evaluation sets. Other datasets like Fake or Real (FoR) [94] and the Chinese Fake Audio Detection (CFAD) [76] offer variations in languages, synthesis methods, or specific conditions. However, those datasets have some limitations: (1) Techniques Coverage Lag: The rapid evolution of voice generation techniques (the arms race described in Section 2) means datasets often lag behind, lacking examples from the latest generators. (2) Aggregation Masking: Traditional evaluations often aggregate diverse fake voices into a single evaluation set. This masks vulnerabilities to specific fake voices types.
Appendix F Fake Voice Evaluation Metrics
Audio was processed at 16kHz where specifically required in PESQ and DNSMOS.
  • TTS Evaluation:
    • Intelligibility: We fed each generated utterance into OpenAI’s Whisper ASR model (medium-en variant) to produce a hypothesis transcript $\hat{T}$ and compare it to the reference transcript TT.[90].
    • Speaker Similarity: For SIM, we used the pretrained WavLM Base+ model [14] to extract deep audio representations. This SIM value is compare with the ground truth same audio clip as a pairing with the fake voice.
    • Perceptual Quality: DNSMOS (overall score) by using official pip package.
    • Spectrum and Signal Fidelity: Peak Signal-to-Noise Ratio (PSNR) quantifies how closely the generated audio’s spectrum matches the reference. We computed the magnitude spectra $X_{\mathrm{ref}}[n]$ and $X_{\mathrm{test}}[n]$ via STFT and normalized both so that the maximum spectral magnitude is 1.
  • Audio Reconstruction Evaluation:
    • Perceptual Quality: PESQ(official), DNSMOS(same as TTS evaluation).
    • Spectrum and Signal Fidelity: SSIM (STFT-based), LS-MSE (log Mel-Spectrograms comparison): we extracted the 80-dimensional mel-spectrogram features using 40 ms Hanning window, 12.5 ms frame shift, 1024-point FFT, and 0 Hz & 12 kHz lower & upper frequency cutoffs following [2], PSNR (same as in TTS evaluation).
  • Computational Complexity:
    • Parameters: We checked the parameters of each models based on their provided checkpoints in their official repository. For Hifi-GAN, MelGAN and ParallelWaveGAN, we used the provided checkpoints from https://github.com/kan-bayashi/ParallelWaveGAN and For FastSpeech2 checkpoints with Hifi-GAN, we used the provided checkpoints in the Google Colab implementation in https://github.com/espnet/espnet.
    • Inference Speed: We use Real-time factor (RTF) to measure the inference speed which can be calculated by the ratio of the time it takes to generate a fake voice to the total duration of this audio, and then aggregated and divided all clips RTFs to directly calculate a mean average value. For text-to-speech generators (including TTS with voice conversion), we include the vocoding time in consideration.
Appendix G Evaluation Metric Normalization
All evaluation metrics are normalized to the open interval $(0,1]$ prior to aggregation in the overall score formulas (Eq. 1–2). Normalization ensures comparability across heterogeneous scales. Let $x$ denote the raw metric value, then the normalized value is given by:
$M_{\text{norm}}=\frac{x-x_{\min}}{x_{\max}-x_{\min}}\cdot(1-\epsilon)+\epsilon,\quad\epsilon>0,$ (1)
where $x_{\min}$ and $x_{\max}$ are the lower and upper bounds for the metric across the evaluation set, and $\epsilon$ is a small constant (e.g., $0.1$) ensuring $M_{\text{norm}}\in(0,1]$. The weights are shown in Table 10.
Appendix H Limitations
  • Model Coverage: Our benchmark evaluates 20+ generators and 8 detector architectures. Despite this breadth, the field moves rapidly—multiple neural codec TTS or generalizable fake voice detector released after 2025 are absent. Consequently, the reported rankings represent a time-stamped evaluation rather than an exhaustive leaderboard.
  • Dataset & Environment: All experiments use clean, single-speaker English downloaded online as original dataset. We do not test noisy or codec audio or low-resource languages whose phonotactics differ markedly. Generalization of both generators and detectors to such conditions therefore remains unverified.
  • ASV Consideration: We focus on fake voice detection countermeasure, but do not measure how each fake voices interferes with automatic speaker verification (ASV) systems. Generator-specific artifacts could degrade ASV scores depending on the back-end. A unified evaluation linking fake voice detectability with ASV system is needed in future work.